### v1

@@ 3.3 Principle 3: Multi-Level Supervision for Latent Representations
Weak Supervision: Noisy or coarse labels, such as position constraints , can be utilized via the trajectory smoothness loss: . Temporal models like Kalman filters [Kalman(1960)] stabilize noisy trajectories in tasks such as autonomous driving. Interval signals as weak supervision can be directly incorporated into the loss function [Mao and Ruchkin(2024)] or combined with contrastive learning to reinforce constraints [Sorokin and Gurevych(2017)].
@@
Combining Supervision Levels: For a given dataset with multiple supervision signals, the training objective must flexibly incorporate relevant terms to ensure that the latent space aligns with the physical world across diverse scenarios. We advocate for using every available supervision opportunity. Given explicit state labels, use direct supervision. When only partial information is available, use weakly supervised constraints to refine the representations. In addition, use self-supervision to extract the sequential physical patterns of observations. Since all these supervisory signals come from the physical world, we expect their multi-level integration to improve physical interpretability.
@@ 3.4 Principle 4: Output Space Partitioning for Verifiability
Ensuring the safety of vision-based autonomy is a critical and open challenge [Teeti et al.(2022)Teeti, Khan, Shahbaz, Bradley, Cuzzolin, and De Raedt, Geng et al.(2024)Geng, Dutta, and Ruchkin]. The verification of such systems remains difficult due to high-dimensional image inputs: traditional techniques struggle to handle this complexity, making it essential to develop new principles for pre-deployment safety guarantees [Althoff(2015), Păsăreanu et al.(2023)Păsăreanu, Mangal, Gopinath, Getir Yaman, Imrie, Calinescu, and Yu]. A recent approach employs a generative image model to map the physical state to the observed image, in turn fed into a state estimator or controller [Katz et al.(2022)Katz, Corso, Strong, and Kochenderfer, Cai et al.(2024)Cai, Fan, and Bak]. Unfortunately, due to uninterpretable latent states, such “verification modulo generative models” does not provide guarantees regarding the physical world. Furthermore, it does not scale to large image sizes.
We propose to partition the generated image into physically meaningful parts to enable the safety verification of vision-based controllers. Specifically, a world model will contain multiple generators of output signals — each dedicated to its own image region. Each generator would be separately verifiable, and the results would be combined to provide world model-wide guarantees. This principle reduces each generator’s size, making the analysis and execution of world models more parallelizable and scalable. Moreover, when applied to physically interpretable latent states, this principle enables the transfer of verification guarantees to the physical world: the generators will represent the relationship between images and physical states, not uninterpretable latent ones.
@@
The question of automatic partitioning of world model outputs can be answered by zero-shot approaches like the Segment Anything Model (SAM) [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.]. Recently, SAM was used to segment images to improve image and safety prediction [Mao et al.(2024a)Mao, Dai, Geng, and Ruchkin]. A similar partitioning was used in the action space to scale up the verification of vision-based controllers via multiple low-dimensional approximations [Geng et al.(2024)Geng, Dutta, and Ruchkin]. Principle 4 propagates the physical meaning from different parts of the world model (established in Principle 1) to its generative outputs, effectively linking the high-dimensional observation space with a lower-dimensional state representation.
@@ 4 Experimental Validation
Principle 2: For the lunar lander, we add a translation to both the observation and position, ensuring the equivariance to translation. For the cart pole, a translation is applied to the cart and a rotation to the pole, with corresponding changes to the latent state. Figures 3A and 3C show that Principle 2 reduces prediction error across all prediction horizons, confirming the value of equivariance.
@@
Principle 4: We partition the original cartpole and lunar lander images into three parts with SAM. Then three smaller decoders are trained to generate each image part, which are combined in the end (as illustrated in Figure 4 in the appendix). The baseline decoder has one linear layer, two convolutional layers, and a 4-dimensional encoded feature map. Our partitioned decoder only contains one linear layer and one smaller convolutional layer. In both the baseline and its partitioned version, the latent space has 4 dimensions for the cartpole and 8 for the lunar lander. The partitioned generator inputs are the exact physical states, while the baseline approach’s uninterpretable latents. Our partitioning reduces 200,259 parameters in the baseline to only 144,665 parameters. Despite that, the reconstruction quality remains comparable, as demonstrated in Table 1.
@@ 5 Future Research Directions
D. Unified Training Pipeline. This paper outlined a variety of training objectives and supervision strategies for physically interpretable world models. However, when their combinations are used, the convergence and stability of training remain elusive [Sener and Koltun(2018)]. We recommend developing an automated training pipeline that will combine and tune different losses to ensure reliable training [Li et al.(2018)Li, Jamieson, DeSalvo, Rostamizadeh, and Talwalkar].
E. Integrating World Models into Classical Autonomy. Physically meaningful states enable high-performance components of world models to serve as state estimators, trajectory predictors, and models for verification [Mao and Ruchkin(2024)]. These capabilities allow combinations between two previously incompatible first-principles and end-to-end learning. We intend to enhance classic autonomy tasks with world-model components to improve their performance while preserving their reliability and verifiability.
@@
Principle 4. Our decoder network maps low-dimensional physical state representations to high-dimensional images using a series of transposed convolutional layers. Using a fully connected layer, the decoder first maps the input state (four-dimensional vector in cartpole; eight-dimensional vector in lunar lander) to a high-dimensional feature space. This produces an intermediate representation of size 3×16×24×24. The image output is further refined through independent transposed convolutional layers, each producing a separate image (three independent layers for each segment image for cartpole and lunar lander). The model is trained using the Adam optimizer with an initial learning rate of 0.001. Training is conducted with mini-batches of size 64, incorporating validation loss tracking to ensure generalization. The loss function is a -weighted combination of the reconstruction MSE of the overall reconstructed image and each segmented part.
For the partitioned loss function in Equation 1, the choice of plays a crucial role in image generation behavior: (a) If is too small ), the model fails to separate the three parts, blending “shadows” of the original image into the outputs; (b) If is too big (), the three parts are completely disconnected, leading to inferior reconstruction quality. Through hyperparameter tuning, we found that setting provides an optimal balance between the quality of the separation and the reconstruction in both case studies.
@@
Table 2 lists the literature on world models and indicates the extent of state/dynamics interpretability and adherence to the proposed principles.
Figure 4 shows example observations and their partitioned reconstructions for Principle 4.
Figure 5 shows the imperfect part-wise reconstruction for inadequate values of .
@@
Short Name Reference Principle 1 Principle 2 Principle 3 Principle 4 State interp. Dyn. interp. WM [Ha and Schmidhuber(2018b)] PlaNet [Hafner et al.(2019a)Hafner, Lillicrap, Fischer, Villegas, Ha, Lee, and Davidson] Weak Weak Dreamer [Hafner et al.(2020)Hafner, Lillicrap, Ba, and Norouzi] Weak Weak G-SWM [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn] Strong Weak Weak Moderate Weak AWM [Kim et al.(2020)Kim, Sano, Freitas, Haber, and Yamins] Weak Weak Weak Plan2Explore [Sekar et al.(2020)Sekar, Rybkin, Daniilidis, Abbeel, Hafner, and Pathak] Weak Weak Weak Pathdreamer [Koh et al.(2021)Koh, Lee, Yang, Baldridge, and Anderson] Weak Weak Strong Weak DreamerV2 [Hafner et al.(2022)Hafner, Lillicrap, Norouzi, and Ba] Weak Weak Weak NSV [Chen et al.(2022)Chen, Huang, Raghupathi, Chandratreya, Du, and Lipson] Strong Strong DayDreamer [Wu et al.(2022)Wu, Escontrela, Hafner, Abbeel, and Goldberg] Weak Weak DreamingV2 [Okada and Taniguchi(2022)] Weak Weak Weak SEN [Park et al.(2022)Park, Biza, Zhao, Meent, and Walters] Strong Moderate STEDI [Nakano et al.(2022)Nakano, Suzuki, and Matsuo] Strong Moderate Strong DriveDreamer [Wang et al.(2023b)Wang, Zhu, Huang, Chen, Zhu, and Lu] Weak Strong Weak GAIA-1 [Hu et al.(2023)Hu, Russell, Yeo, Murez, Fedoseev, Kendall, Shotton, and Corrado] Strong Moderate IFactor [Liu et al.(2023)Liu, Xiao, and Belta] Moderate Moderate IRIS [Micheli et al.(2023a)Micheli, Alonso, and Fleuret] Weak Weak MTS3 [Shaj et al.()Shaj, Zadeh, Demir, Douat, and Neumann] Weak Weak Weak Weak Denoised MDP [Wang et al.(2023a)Wang, Du, Torralba, Isola, Zhang, and Tian] Moderate Moderate WM2WM [Wong et al.(2023)Wong, Grand, Lew, Goodman, Mansinghka, Andreas, and Tenenbaum] Strong Moderate MWM [Seo et al.(2023)Seo, Hafner, Liu, Liu, James, Lee, and Abbeel] Weak Weak OccWorld [Zheng et al.(2023)Zheng, Chen, Huang, Zhang, Duan, and Lu] Strong Strong Strong RAP [Hao et al.(2023)Hao, Gu, Ma, Hong, Wang, Wang, and Hu] Moderate S4WM [Deng et al.(2023)Deng, Park, and Ahn] Moderate SWIM [Mendonca et al.(2023)Mendonca, Bahl, and Pathak] Moderate Moderate Moderate TWM [Robine et al.(2023)Robine, Höftmann, Uelwer, and Harmeling] Weak Weak UniWorld [Min et al.(2023)Min, Zhao, Xiao, Nie, and Dai] Strong Strong Strong WorldCloner [Balloch et al.(2023)Balloch, Lin, Peng, Hussain, Srinivas, Wright, Kim, and Riedl] Strong Moderate Strong THICK [Gumbsch et al.(2023)Gumbsch, Sajid, Martius, and Butz] Weak Weak Weak AVID [Rigter et al.(2024)Rigter, Gupta, Hilmkil, and Ma] Strong Moderate CMIL [Kolev et al.(2024)Kolev, Rafailov, Hatch, Wu, and Finn] Strong Weak DreamerV3 [Hafner et al.(2024)Hafner, Pasukonis, Ba, and Lillicrap] Weak Weak Weak Weak DriveWorld [Min et al.(2024)Min, Zhao, Xiao, Zhao, Xu, Zhu, Jin, Li, Guo, Xing, Jing, Nie, and Dai] Strong Moderate Weak DWM [Ding et al.(2024)Ding, Zhang, Tian, and Zheng] GaussianWorld [Zuo et al.(2024)Zuo, Zheng, Huang, Zhou, and Lu] Moderate Strong Strong Moderate Genie [Bruce et al.(2024)Bruce, Dennis, Edwards, Parker-Holder, Shi, Hughes, Lai, Mavalankar, Steigerwald, Apps, Aytar, Bechtle, Behbahani, Chan, Heess, Gonzalez, Osindero, Ozair, Reed, Zhang, Zolna, Clune, Freitas, Singh, and Rocktäschel] Moderate Moderate HarmonyWM [Ma et al.(2024)Ma, Wu, Feng, Xiao, Li, Hao, Wang, and Long] Weak Weak OccWM [Zhang et al.(2024)Zhang, Xue, Yan, Zhang, Qiu, Bai, Liu, Cui, and Li] Moderate Strong Strong Weak CovWM [Popov et al.(2024)Popov, Degirmenci, Wehr, Hegde, Oldja, Kamenev, Douillard, Nistér, Muller, Bhargava, Birchfield, and Smolyanskiy] Weak Weak Weak Weak NWM [Bar et al.(2024)Bar, Zhou, Tran, Darrell, and LeCun] OccLLaMA [Wei et al.(2024)Wei, Yuan, Li, Hu, Gan, and Ding] Strong Strong PIWM [Mao and Ruchkin(2024)] Strong Strong Moderate Strong R-AIF [Nguyen et al.(2024)Nguyen, Yang, Buckley, and Ororbia] Weak Weak Weak RenderWorld [Yan et al.(2024)Yan, Dong, Shao, Lu, Haiyang, Liu, Wang, Wang, Wang, Remondino, and Ma] Strong Strong Strong Think2Drive [Li et al.(2024)Li, Jia, Wang, and Yan] Weak Weak TransDreamer [Chen et al.(2024)Chen, Wu, Yoon, and Ahn] Weak Weak Weak VisualPredicator [Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis] Strong Strong Moderate WorldGPT [Ge et al.(2024)Ge, Huang, Zhou, Li, Wang, Tang, and Zhuang] Moderate Moderate Moderate Moderate Our future vision Strong Strong Strong Strong Strong Strong

### v2

@@ 3.3 Principle 3: Multi-Level and Multi-Strength Supervision for Latent Representations
Weak Supervision: Noisy or coarse labels, such as position constraints , can be utilized via the trajectory smoothness loss: . Temporal models like Kalman filters [Kalman(1960)] stabilize noisy trajectories in tasks such as autonomous driving. Interval signals as weak supervision can be directly incorporated into the loss [Mao and Ruchkin(2024)] or combined with contrastive learning to reinforce constraints [Sorokin and Gurevych(2017)].
@@
Combining Supervision Levels: For a given dataset with supervision signals (e.g., full trajectories, state variables, or interval constraints), the training objective should integrate matching losses to align the latent space with physical semantics. We advocate for using every available supervision. Given explicit state labels, use direct supervision. When only partial information is available, use weakly supervised constraints to refine the representations.
@@ 3.4 Principle 4: Output Space Partitioning for Verifiability
Ensuring the safety of vision-based autonomy is a critical and open challenge [Teeti et al.(2022)Teeti, Khan, Shahbaz, Bradley, Cuzzolin, and De Raedt, Geng et al.(2024)Geng, Dutta, and Ruchkin, Fremont et al.(2019)Fremont, Dreossi, Ghosh, Yue, Sangiovanni-Vincentelli, and Seshia]. The verification of such systems is difficult due to high-dimensional image inputs: classical techniques cannot handle this complexity, motivating new principles for pre-deployment safety guarantees [Althoff(2015), Păsăreanu et al.(2023)Păsăreanu, Mangal, Gopinath, Getir Yaman, Imrie, Calinescu, and Yu]. One such attempt is to employ a generative image model to overapproximate observations from a given physical state and feed them into a state estimator or controller [Katz et al.(2022)Katz, Corso, Strong, and Kochenderfer, Cai et al.(2024)Cai, Fan, and Bak]. Sadly, due to uninterpretable latent states, such “verification modulo generative models” does not provide guarantees for the physical world. Furthermore, decoder verification does not scale to large images.
To reduce decoder complexity, we propose to partition the generated image into physically meaningful parts. Specifically, a world model will contain multiple generators of output signals — each dedicated to its own object in the image. Each generator would be separately verifiable, and the results would be combined to provide world model-wide guarantees. This principle reduces each generator’s size, making the verification of such world models tractable.
When applied to physically interpretable latent states, this principle can transfer verification guarantees to the physical world: the generators would represent the relationship between images and physical states, not uninterpretable latent ones. Specifications are written at the level of interpretable states (rather than images), after the information has been propagated through perception, control, and dynamics. This process grounds the verification in physical states rather than images.
@@
The question of automatic partitioning of world model outputs can be answered by zero-shot approaches like the Segment Anything Model (SAM) [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.]. Recently, SAM was used to segment images to improve image and safety prediction [Mao et al.(2024a)Mao, Dai, Geng, and Ruchkin]. A similar partitioning was used in the action space to scale up the verification of vision-based controllers via multiple low-dimensional approximations [Geng et al.(2024)Geng, Dutta, and Ruchkin]. Principle 4 propagates the physical meaning from different parts of the world model (established in Principle 1) to its generative outputs, effectively linking the high-dimensional observation with a lower-dimensional representation.
This principle has two remaining limitations. First, as the number of objects increases, partitioning becomes increasingly difficult, as in autonomous driving tasks with dynamic objects like cars and pedestrians. Additionally, the gap between a world model and the real world still needs to be formally quantified to obtain guarantees, which is an open problem for future research.
@@ 4 Experimental Validation
Principle 2: We specify a function that shifts the lunar lander’s position, and a corresponding function that shifts the latent state. For cart pole, we shift both the rotation and position, with corresponding changes made to the latent state. Figures 3A and 3C show that Principle 2 reduces prediction error across all prediction horizons, confirming the value of equivariance. While this principle improves performance on lunar lander, it has less of an effect on the cart pole. We hypothesize that this principle benefits more complex and partially observable systems.
@@
Principle 4: We partition the original cartpole and lunar lander images into three parts with SAM, training three smaller decoders for each and combining them as shown in Figure 4 in the Appendix. The partitioned generator inputs are the exact physical states, while the baseline approach’s uninterpretable latents. Our partitioning reduces the baseline’s parameters by 27.7% while keeping a comparable reconstruction quality remains comparable, as per Table 1. Though based on simple environments, these standard benchmarks help isolate each principle’s effect. We plan to extend validation to more complex domains like 3D navigation and visual robotic manipulation.
@@ 5 Future Research Directions
D. Unified Training Pipeline. We outlined several training objectives and supervision strategies for world models. However, when their combinations are used, the convergence and stability of training remain elusive [Sener and Koltun(2018)]. We recommend developing an automated training pipeline that will combine and tune different losses to ensure reliable training [Li et al.(2018)Li, Jamieson, DeSalvo, Rostamizadeh, and Talwalkar].
E. Integrating World Models into Classical Autonomy. Physically meaningful states enable high-performance components of world models to serve as state estimators, trajectory predictors, and verification models [Mao and Ruchkin(2024)]. This allows combining the previously incompatible first-principles and end-to-end learning models. We intend to improve the performance of classic autonomy tasks with world-model components while preserving their reliability and verifiability.
@@
Principle 4. Our decoder network maps low-dimensional physical state representations to high-dimensional images using a series of transposed convolutional layers. The baseline decoder has one linear layer, two convolutional layers, and a 4-dimensional encoded feature map. Our partitioned decoder only contains one linear layer and one smaller convolutional layer. Using a fully connected layer, the decoder first maps the input state (four-dimensional vector in cartpole; eight-dimensional vector in lunar lander) to a high-dimensional feature space. This produces an intermediate representation of size 3×16×24×24. The image output is further refined through independent transposed convolutional layers, each producing a separate image (three independent layers for each segment image for cartpole and lunar lander). The model is trained using the Adam optimizer with an initial learning rate of 0.001. Training is conducted with mini-batches of size 64, incorporating validation loss tracking to ensure generalization. The loss function is a -weighted combination of the reconstruction MSE of the overall reconstructed image and each segmented part.
For the partitioned loss function in Definition 3.4, the choice of plays a crucial role in image generation behavior: (a) If is too small ), the model fails to separate the three parts, blending “shadows” of the original image into the outputs; (b) If is too big (), the three parts are completely disconnected, leading to inferior reconstruction quality. Through hyperparameter tuning, we found that setting provides an optimal balance between the quality of the separation and the reconstruction in both case studies.
@@
Short Name Reference Principle 1 Principle 2 Principle 3 Principle 4 State interp. Dyn. interp. WM [Ha and Schmidhuber(2018)] PlaNet [Hafner et al.(2019)Hafner, Lillicrap, Fischer, Villegas, Ha, Lee, and Davidson] Weak Weak Dreamer [Hafner et al.(2020)Hafner, Lillicrap, Ba, and Norouzi] Weak Weak G-SWM [Lin et al.(2020)Lin, Wu, Peri, Fu, Jiang, and Ahn] Strong Weak Weak Moderate Weak AWM [Kim et al.(2020)Kim, Sano, Freitas, Haber, and Yamins] Weak Weak Weak Plan2Explore [Sekar et al.(2020)Sekar, Rybkin, Daniilidis, Abbeel, Hafner, and Pathak] Weak Weak Weak Pathdreamer [Koh et al.(2021)Koh, Lee, Yang, Baldridge, and Anderson] Weak Weak Strong Weak DreamerV2 [Hafner et al.(2022)Hafner, Lillicrap, Norouzi, and Ba] Weak Weak Weak NSV [Chen et al.(2022)Chen, Huang, Raghupathi, Chandratreya, Du, and Lipson] Strong Strong DayDreamer [Wu et al.(2022)Wu, Escontrela, Hafner, Abbeel, and Goldberg] Weak Weak DreamingV2 [Okada and Taniguchi(2022)] Weak Weak Weak SEN [Park et al.(2022)Park, Biza, Zhao, Meent, and Walters] Strong Moderate STEDI [Nakano et al.(2022)Nakano, Suzuki, and Matsuo] Strong Moderate Strong DriveDreamer [Wang et al.(2023b)Wang, Zhu, Huang, Chen, Zhu, and Lu] Weak Strong Weak GAIA-1 [Hu et al.(2023)Hu, Russell, Yeo, Murez, Fedoseev, Kendall, Shotton, and Corrado] Strong Moderate IFactor [Liu et al.(2023)Liu, Xiao, and Belta] Moderate Moderate IRIS [Micheli et al.(2023)Micheli, Alonso, and Fleuret] Weak Weak MTS3 [Shaj et al.()Shaj, Zadeh, Demir, Douat, and Neumann] Weak Weak Weak Weak Denoised MDP [Wang et al.(2023a)Wang, Du, Torralba, Isola, Zhang, and Tian] Moderate Moderate WM2WM [Wong et al.(2023)Wong, Grand, Lew, Goodman, Mansinghka, Andreas, and Tenenbaum] Strong Moderate MWM [Seo et al.(2023)Seo, Hafner, Liu, Liu, James, Lee, and Abbeel] Weak Weak OccWorld [Zheng et al.(2023)Zheng, Chen, Huang, Zhang, Duan, and Lu] Strong Strong Strong RAP [Hao et al.(2023)Hao, Gu, Ma, Hong, Wang, Wang, and Hu] Moderate S4WM [Deng et al.(2023)Deng, Park, and Ahn] Moderate SWIM [Mendonca et al.(2023)Mendonca, Bahl, and Pathak] Moderate Moderate Moderate TWM [Robine et al.(2023)Robine, Höftmann, Uelwer, and Harmeling] Weak Weak UniWorld [Min et al.(2023)Min, Zhao, Xiao, Nie, and Dai] Strong Strong Strong WorldCloner [Balloch et al.(2023)Balloch, Lin, Peng, Hussain, Srinivas, Wright, Kim, and Riedl] Strong Moderate Strong THICK [Gumbsch et al.(2023)Gumbsch, Sajid, Martius, and Butz] Weak Weak Weak AVID [Rigter et al.(2024)Rigter, Gupta, Hilmkil, and Ma] Strong Moderate CMIL [Kolev et al.(2024)Kolev, Rafailov, Hatch, Wu, and Finn] Strong Weak DreamerV3 [Hafner et al.(2024)Hafner, Pasukonis, Ba, and Lillicrap] Weak Weak Weak Weak DriveWorld [Min et al.(2024)Min, Zhao, Xiao, Zhao, Xu, Zhu, Jin, Li, Guo, Xing, Jing, Nie, and Dai] Strong Moderate Weak DWM [Ding et al.(2024)Ding, Zhang, Tian, and Zheng] GaussianWorld [Zuo et al.(2024)Zuo, Zheng, Huang, Zhou, and Lu] Moderate Strong Strong Moderate Genie [Bruce et al.(2024)Bruce, Dennis, Edwards, Parker-Holder, Shi, Hughes, Lai, Mavalankar, Steigerwald, Apps, Aytar, Bechtle, Behbahani, Chan, Heess, Gonzalez, Osindero, Ozair, Reed, Zhang, Zolna, Clune, Freitas, Singh, and Rocktäschel] Moderate Moderate HarmonyWM [Ma et al.(2024)Ma, Wu, Feng, Xiao, Li, Hao, Wang, and Long] Weak Weak OccWM [Zhang et al.(2024)Zhang, Xue, Yan, Zhang, Qiu, Bai, Liu, Cui, and Li] Moderate Strong Strong Weak CovWM [Popov et al.(2024)Popov, Degirmenci, Wehr, Hegde, Oldja, Kamenev, Douillard, Nistér, Muller, Bhargava, Birchfield, and Smolyanskiy] Weak Weak Weak Weak NWM [Bar et al.(2024)Bar, Zhou, Tran, Darrell, and LeCun] OccLLaMA [Wei et al.(2024)Wei, Yuan, Li, Hu, Gan, and Ding] Strong Strong PIWM [Mao and Ruchkin(2024)] Strong Strong Moderate Strong R-AIF [Nguyen et al.(2024)Nguyen, Yang, Buckley, and Ororbia] Weak Weak Weak RenderWorld [Yan et al.(2024)Yan, Dong, Shao, Lu, Haiyang, Liu, Wang, Wang, Wang, Remondino, and Ma] Strong Strong Strong Think2Drive [Li et al.(2024)Li, Jia, Wang, and Yan] Weak Weak TransDreamer [Chen et al.(2024)Chen, Wu, Yoon, and Ahn] Weak Weak Weak VisualPredicator [Liang et al.(2024)Liang, Kumar, Tang, Weller, Tenenbaum, Silver, Henriques, and Ellis] Strong Strong Moderate WorldGPT [Ge et al.(2024)Ge, Huang, Zhou, Li, Wang, Tang, and Zhuang] Moderate Moderate Moderate Moderate Our future vision Strong Strong Strong Strong Strong Strong